OpenStudy (anonymous):

5 divided by 6 minus 3 dived by 8

OpenStudy (anonymous):

5/6-3/8=20-9/24=11/24

sammixboo (sammixboo):

\(\sf\color{magenta}{5 ÷ 6 - 3 ÷ 8}\) First you have to divide first so what is \(\sf\color{magenta}{5 ÷ 6}\) and \(\sf\color{magenta}{ 3 ÷ 8}\) One you have divided those two problems you must take the answers to the division problems and subtract them together. Once you have done that is your final answer :)

sammixboo (sammixboo):

Make sure you tell us what your final answer is so we can see if you done it right or wrong \(\large\sf\color{magenta}{:)}\)

sammixboo (sammixboo):

@bjohns978